OpenStudy (anonymous):

Given A = [ 3 -2 -6 4 12 -8 ], find one nontrivial solution Ax = 0 by inspection.

OpenStudy (anonymous):

More specifically, what is meant by inspection?

OpenStudy (anonymous):

by inspection just means just by looking at it

OpenStudy (anonymous):

I'm not sure what exactly to look for, but I know that there's going to be infinitely many solutions (there's more rows than columns, so there's gonna be a row without pivots).

OpenStudy (anonymous):

...wait, does that mean that since there's not a pivot in every column, that I could just throw any solution in and it would work?

OpenStudy (anonymous):

every row*

OpenStudy (anonymous):

i think the matrix you have shown is incorrect have you considered x=[0; 0] yet?

OpenStudy (anonymous):

But I'm looking for a nontrivial solution.

OpenStudy (anonymous):

are you sure you're not missing a negative sign on the top row?

OpenStudy (anonymous):

Whoops! You're right.

OpenStudy (anonymous):

...waitaminute.

OpenStudy (anonymous):

The second two rows can be reduced to zero rows.

OpenStudy (anonymous):

if you look at the first row, then look at the last 2 rows then realize that the last 2 rows are multiples of the 1st row then solve for just the first row

OpenStudy (anonymous):

So now we've got [ 3 -2; 0 0; 0 0;].

OpenStudy (anonymous):

But I feel like I'm still forgetting a step.

OpenStudy (anonymous):

I don't think x = [3 -2] is the solution.

OpenStudy (anonymous):

x=[x;y] now solve for x and y for Ax=0

OpenStudy (anonymous):

\[x=\left(\begin{matrix}x_1 \\ x_2\end{matrix}\right)\] \[(3 -2)\left(\begin{matrix}x_1 \\ x_2\end{matrix}\right)=0\]

OpenStudy (anonymous):

So x1 and x2 would have to be zero, then?

OpenStudy (anonymous):

you said nontrivial

OpenStudy (anonymous):

basically make up a number for x1, then solve for x2 then rewrite x

OpenStudy (anonymous):

Ok, give me a second here...

OpenStudy (anonymous):

So I could set x1 = 2 and x2 = 3, which would mean the result would be (3*2)+(3*-2)=6-6=0, which means the solution would be x = [2 3]?

OpenStudy (anonymous):

and also [4 6] and [8 12] etc etc etc

OpenStudy (anonymous):

awesome, thanks for the help!